OpenStudy (anonymous):

How many solutions exist for the system y = 2x + 5 and y = -0.2x - 4? A. no solution B. one solution C. infinitely many solutions

OpenStudy (amistre64):

if we scale the 2nd by -4 we get 2x + 16

OpenStudy (amistre64):

same x parts different constants; no solutions

OpenStudy (amistre64):

opps, read .5 instead of .2

OpenStudy (amistre64):

same concept tho :)

OpenStudy (amistre64):

y = 2x + 5 -y = 0.2x + 4 ------------- 0 = 2.2x + 9 x = -9/2.2

OpenStudy (amistre64):

if that x produces the same ys were good

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y+%3D+2x+%2B+5+and+y+%3D+-0.2x+-+4 says were good with one solution

OpenStudy (anonymous):

Change -0.2 to -(2/10) and solve the following for x and y:\[\left\{y=2 x+5,y=-\frac{x}{5}-4\right\} \]\[\left\{x=-\frac{45}{11},y=-\frac{35}{11}\right\} \]